# On the feasibility of performing quantum chemistry calculations on quantum computers

### v3

@@ Abstract
where $c^{\dagger}_{i}$ (resp. $c_{i}$) creates (resp. destroys) an electron on the spin-orbital $i\in\{1,...,M\}$, $h_{ij}$ corresponds to the kinetic and nuclei Hamiltonian matrix elements and $v_{ijkl}$ corresponds to the electron-electron Coulomb matrix elements. The goal of VQE and QPE is to find the ground state $|\Psi_{0}\rangle$ of $H$ and in particular calculate the associated energy $E_{0}$.
Such a problem can be addressed by a large spectrum of classical techniques with different precisions and computational costs. Mean-field techniques (improved with e.g. GW diagrammatic approaches) can reach a precision of $\approx 10$  mHa for up to $M>5000$ and hundreds of electrons (see e.g. [18] for typical calculations), which is sufficient for calculating dissociation energies (including phonon spectra), the energy liberated by a chemical reaction, or ionization energies. Reaching an accuracy of ${\approx~{}1\mathrm{mHa}~{}\approx~{}300\mathrm{K}}$, needed to compute activation energies, requires more precise techniques such as coupled cluster techniques, whose computational complexity scale as $N^{2}(M-N)^{4}$ (up to two electron-hole excitations), $N^{3}(M-N)^{5}$ (three excitations) or more, making these approaches impractical beyond a few hundreds orbitals [19]. The above approaches work well for many molecules. However they typically fail for strongly correlated molecules (e.g. containing transition metals), which require even more complex methods such as the density-matrix renormalization group method (DMRG, [20]). These methods can typically handle only tens of orbitals. It is this latter case that has been identified as the early target for quantum algorithms, in particular in the context of nitrogen fixation for fertilizers [21].
Before introducing the two criteria, let us emphasize that the size $M$ of the basis set plays a very important role in the very definition of the quality of a ground-state energy estimation, as pointed out in Ref. [12]. One indeed needs to distinguish precision—the ability to correctly solve the many-body problem for given $N$ and $M$—and accuracy—the ability to reach large enough $M$ to approach the complete basis set limit $M\rightarrow\infty$ for a given $N$. For instance, the ”spherical cow” of the VQE algorithm, namely the benchmark of the H2 energy versus H--H distance, almost always uses only a single orbital per hydrogen atom (STO-3G or STO-6G basis set) and corresponds to an accuracy of at best $28$mHa, even if one reaches perfect precision. The true target of quantum computers is chemical accuracy ($<1.6$mHa), which always means using much larger basis sets. This means more qubits and, as we shall see, more severe requirements on the hardware.
@@ I A criterion for the variational quantum eigensolver (VQE): maximum hardware error
We start our analysis with VQE, which has been proposed as a possible algorithm that could be usable with noisy quantum computers. VQE suffers from three important difficulties: (I) its sensitivity to statistical uncertainty, (II) the difficulty to perform the optimization of the wavefunction and (III) its sensitivity to hardware imprecisions and decoherence. In this section we briefly review the VQE algorithm as well as problems I and II, then proceed with an in-depth discussion of problem III. We conclude with a point-by-point comparison of VQE and its natural classical counterpart, the variational Monte-Carlo (VMC) algorithm.
@@ I.1 Description of VQE and its challenges
with classical optimization algorithms. The use of a quantum computer to prepare the variational wavefunction allows to consider classes of wavefunctions that could a priori not be accessed using classical computers. Starting from a given problem, one constructs the corresponding Hamiltonian representation, choosing a given molecular basis (e.g. Gaussian basis functions) and representation (e.g. second quantization). Then, the Hamiltonian is encoded by e.g. a Jordan-Wigner transformation into a qubit representation. One key point of VQE is the choice of an ansatz for the variational wavefunction, namely the parameterized quantum circuit $U(\vec{\theta})$. The ansatz must be ”expressive” enough so that it covers the interesting part of the Hilbert space, while using as few parameters as possible, and allowing for a shallow circuit representation due to the limited circuit depths and run times available in NISQ, real-life experiments.
@@
with $H_{\alpha}=\sum_{k_{\alpha}}\lambda_{k_{\alpha}}P_{k_{\alpha}}$, with $P_{k_{\alpha}}$ of the form $\sigma_{1}\otimes\cdots\sigma_{n}$ (where $\sigma_{i}$ is one of the Pauli matrices acting on qubit $i$ ($X$, $Y$, $Z$ or identity)). The expectation value of each term $H_{\alpha}$ is measured separately: for each $\alpha$, one repeatedly prepares state $|\Psi_{V}\rangle$ and measures each $P_{k_{\alpha}}$ in the common eigenbasis of the $P_{k_{\alpha}}$’s. The outcomes are averaged to get an estimate of $\langle\Psi_{V}|P_{k_{\alpha}}|\Psi_{V}\rangle$. The statistical error stemming from the finite number $N_{s}$ of repetitions, or shots, is of the form $\epsilon\approx\sigma(P_{k_{\alpha}})/\sqrt{N_{s}}$ where $\sigma(P_{k_{\alpha}})$ is the standard deviation of individual measurements. Therefore, the run time of computing $E_{V}$, which is proportional to $N_{s}$, scales as $1/\epsilon^{2}$. This scaling is a direct consequence of the fact that energy measurements in VQE are based on a classical averaging of individual measurement outcomes. Even more advanced methods like classical shadows [22] also suffer from this scaling, although they can alleviate the dependence of the run time on the number of terms to be measured.
This unfavorable scaling of VQE leads to quite large run time estimates if good enough accuracies $\epsilon$ are to be reached: in 2015, a study estimated that to evaluate the ground state energy of Fe2Se2 in the minimal STO33-3- 3g orbital basis with chemical precision would require a prohibitive number of $10^{19}$ samples [15]. Another study estimated the runtime to estimate the combustion energy for methane to be of 1.9 days and 71 days for ethanol, finding no practical advantage compared to state-of-the-art classical techniques (see Table II in [14]). In a recent review, combining the state of the art methods in each part of the VQE pipeline and assuming a gate time of $100$ ns, the total runtime for one iteration of VQE on the Cr2 molecule is estimated at 25 days. Taking into account the number of iterations required to reach the minimum in parameter space yields an estimate of 24 years [10].
@@ I.2 The decoherence versus precision criterion
In this section, we disregard statistical uncertainty and barren plateaus, and aim at answering a more pragmatic short-term question: what is the level of decoherence sustainable to attain a targeted precision on the ground state energy? (Problem III). Note that we use the terminology ”decoherence” generically here to account for all the imperfections of the quantum hardware including actual decoherence, but also more mundane limitations to the precisions such as the effect of crosstalk, 1/f noise, imprecise pulse sequence…[27].
@@ I.2.1 Derivation of the criterion
We measure the effect of the noise on the hardware with the fidelity $F=\langle\Psi_{V}|\rho|\Psi_{V}\rangle$. It expresses how the density matrix $\rho$ of the quantum computer after a noisy execution of the variational circuit $U(\vec{\theta})$ differs from the expected one, $|\Psi_{V}\rangle\langle\Psi_{V}|$. A $F<1$ implies that $\rho$ can always be written in the form
@@
with $E_{\rm noise}={\rm Tr}(\rho_{\rm noise}H)$. A large corpus of experiments and theory [28], including the seminal ”quantum supremacy” experiment by Google [29], shows that the fidelity decays exponentially with the total number of applied gates $N_{g}$,
@@
where $\epsilon$ is the average error per gate. In the leading quantum hardware, this error is dominated by the two-qubit gates for which $\epsilon\leq 1\%$. The fact that different errors get combined in a multiplicative way can be put on firm mathematical basis only for certain error models, but is widely expected to hold more generally.
Suppose that one targets a precision $\eta_{\rm chem}$ in the calculation with $\eta_{\rm chem}\ll E_{\mathrm{noise}}-E_{V}$. It follows that $F$ must be very close to unity. Combining the above equations, imposing $\Delta E\leq\eta_{\rm chem}$ leads to
$\epsilon\leq\frac{\eta_{\rm chem}}{(E_{\rm noise}-E_{V})N_{g}}.$ (8)
This is our quantitative criterion for using VQE on a given hardware, molecule and ansatz. Let us emphasize that, in practice, $N_{g}$ depends implicitly on $\eta_{\rm chem}$: good precision requires a complex ansatz, and hence a large number of gates. Note that $\epsilon$ is the effective error that one obtains after one has done one’s best to decrease the error level, including any quantum error correction scheme and/or error mitigation (which is discussed below).
Figure 1: Schematic of the difference between the hardware spectrum and that of the studied molecule. The VQE ansatz is close to the ground state of the molecule but is made of arbitrary high energy states of the hardware. Conversely, the hardware noise will populate arbitrary high excited states of the studied molecule.
@@ I.2.2 Discussion: scaling of the $(E_{\mathrm{noise}}-E_{V})$ energy scale
We now argue that the energy scale in the denominator, $(E_{\mathrm{noise}}-E_{V})$, is generically a large energy scale—of the order of the bare matrix elements of the Hamiltonian, Hartrees or tens of Hartrees—and even generically scales very unfavorably as the square of the number of electrons.
To understand the origin of this behaviour, we need to remember that there are two very different Hamiltonians in the problem with two very different energy spectra: the target Hamiltonian $H$ (the one we want to find the ground state of) and the Hamiltonian $H^{\rm hw}$ that describes the quantum hardware (sometimes called the ”resource Hamiltonian”, see e.g [30]). More generally, the dynamics of the quantum computer hardware is described by a time-dependent Lindbladian, but for the purpose of this discussion, it is sufficient to consider its Hamiltonian. We thus have two sets of eigenstates and eigenenergies,
@@
Our main observation is that in general the target Hamiltonian $H$ is very different from the hardware Hamiltonian $H^{\rm hw}$. Indeed the same hardware Hamiltonian must be used for calculating different molecules so it generically has no correlations with the latter. This is a constraint a gate-based quantum circuit approach (as opposed to an analog quantum simulation) must obey in exchange for its universality. This situation is different from e.g annealing algorithms where the hardware Hamiltonian is supposed to be slowly deformed into the target Hamiltonian to end up in a low-energy state of the target Hamiltonian. This is also different from analog computing models, which strive to construct hardware Hamiltonians that are as close as possible to the target Hamiltonian, and may thus have similar low-energy subspaces. In other words, gate-based models, to gain universality, lose the physical constraints (for instance, symmetries) that could guarantee that the hardware Hamiltonian indeed shares properties with the target Hamiltonian.
@@
with $c_{i}$ having non zero components in the entire spectrum $E_{i}^{\rm hw}$ (see Fig. 1 for a schematic). Conversely the hardware eigenstates involve very high energies of the molecule (possibly even states with different number of electrons). Why is this observation important? Because the mechanisms that lead to imprecisions and decoherence usually have a simple structure related to the hardware spectrum. This in turns means that these mechanisms will be structureless with respect to the target spectrum. A simple example is energy relaxation of the hardware—a ubiquitous process. Its fixed point is the thermal density matrix $\rho\propto\sum_{i}e^{-E_{i}/kT}|\Psi^{\rm hw}_{i}\rangle\langle\Psi^{\rm hw}_{i}|$. As another example, consider the admixture of a small amount $(1-F)$ of, say the ground state, $|\Psi^{\rm hw}_{0}\rangle\langle\Psi^{\rm hw}_{0}|$ in the density matrix: it will lead to a noise energy $E_{\rm noise}=\langle\Psi^{\rm hw}_{0}|H|\Psi^{\rm hw}_{0}\rangle$. Since $|\Psi^{\rm hw}_{0}\rangle$ is composed of a superposition of ”target” states $|\Psi_{i}\rangle$ of arbitrary ”target” energies $E_{i}$, it follows that $E_{\rm noise}$ typically lies in the middle of the spectrum of the target Hamiltonian,
@@
where the magnitude of the constants $a$ and $b$ will depend on the details of the problem and of the noise model. $a$ and $b$ are large energy scales, of the same order as the matrix elements $h_{ij}$ and $v_{ijkl}$ (from Hartrees to tens of Hartrees, depending on the basis sets).

### v4

# Feasibility of performing quantum chemistry calculations on quantum computers

@@ Abstract
where $c^{\dagger}_{i}$ (respectively $c_{i}$) creates (respectively destroys) an electron on the spin-orbital $i\in\{1,...,M\}$, $h_{ij}$ corresponds to the kinetic and nuclei Hamiltonian matrix elements and $v_{ijkl}$ corresponds to the electron-electron Coulomb matrix elements. The goal of VQE and QPE is to find the ground state $|\Psi_{0}\rangle$ of $H$ and in particular calculate the associated energy $E_{0}$.
Such a problem can be addressed by a large spectrum of classical techniques with different precisions and computational costs. Mean-field techniques (improved with e.g. GW diagrammatic approaches) can reach a precision of $\approx 10$  mHa for up to $M>5000$ and hundreds of electrons (see e.g. [18] for typical calculations), which is sufficient for calculating dissociation energies (including phonon spectra), the energy liberated by a chemical reaction, or ionization energies. Reaching an accuracy of $\approx~1\mathrm{~mHa}~\approx~300\mathrm{~K}$}, needed to compute activation energies, requires more precise techniques such as coupled cluster techniques, whose computational complexity scales as $N^{2}(M-N)^{4}$ (up to two electron-hole excitations), $N^{3}(M-N)^{5}$ (three excitations) or more, making these approaches impractical beyond a few hundred orbitals [19]. The above approaches work well for many molecules. However they typically fail for strongly correlated molecules (e.g. containing transition metals), which require even more complex methods such as the density-matrix renormalization group method (DMRG, [20]). These methods can typically handle only tens of orbitals. It is this latter case that has been identified as the early target for quantum algorithms, in particular in the context of nitrogen fixation for fertilizers [21], although addressing this problem might not be as ground breaking as initially surmised [5].
Before introducing the two criteria, let us emphasize that the size $M$ of the basis set plays a very important role in the very definition of the quality of a ground-state energy estimation, as pointed out in Ref. [12]. One indeed needs to distinguish precision—the ability to correctly solve the many-body problem for given $N$ and MM—and accuracy—the ability to reach large enough $M$ to approach the complete basis set limit $M\rightarrow\infty$ for a given $N$. For instance, the ”spherical cow” of the VQE algorithm, namely the benchmark of the H2 energy versus H-H distance, almost always uses only a single orbital per hydrogen atom (STO-3G or STO-6G basis set) and corresponds to an accuracy of at best $28$ mHa, even if one reaches perfect precision. The true target of quantum computers is chemical accuracy (<1.6<1.6 mHa), which always means using much larger basis sets. This means more qubits and, as we shall see, more severe requirements on the hardware.
@@ I A criterion for the variational quantum eigensolver (VQE): maximum hardware error
We start our analysis with VQE, which has been proposed as a possible algorithm that could be usable with noisy quantum computers. VQE suffers from three important difficulties: (I) its sensitivity to statistical uncertainty, (II) the difficulty to perform the optimization of the wavefunction and (III) its sensitivity to hardware imprecisions and decoherence. In this section we briefly review the VQE algorithm as well as problems I and II, then proceed with an in-depth discussion of problem III. We conclude with a point-by-point comparison of VQE and its natural classical counterpart, the variational Monte Carlo (VMC) algorithm.
@@ I.1 Description of VQE and its challenges
with classical optimization algorithms. The use of a quantum computer to prepare the variational wavefunction allows to consider classes of wavefunctions that could a priori not be accessed using classical computers. Starting from a given problem, one constructs the corresponding Hamiltonian representation, choosing a given molecular basis (e.g., Gaussian basis functions) and representation (e.g., second quantization). Then, the Hamiltonian is encoded by e.g., a Jordan-Wigner transformation into a qubit representation. One key point of VQE is the choice of an ansatz for the variational wavefunction, namely the parameterized quantum circuit $U(\vec{\theta})$. The ansatz must be ”expressive” enough so that it covers the interesting part of the Hilbert space, while using as few parameters as possible, and allowing for a shallow circuit representation due to the limited circuit depths and run times available in NISQ, real-life experiments.
@@
with $H_{\alpha}=\sum_{k_{\alpha}}\lambda_{k_{\alpha}}P_{k_{\alpha}}$, with $P_{k_{\alpha}}$ of the form $\sigma_{1}\otimes\cdots\otimes\sigma_{n}$ (where $\sigma_{i}$ is one of the Pauli matrices acting on qubit $i$ ($X$, $Y$, $Z$ or identity)). The expectation value of each term $H_{\alpha}$ is measured separately: for each $\alpha$, one repeatedly prepares state $|\Psi_{V}\rangle$ and measures each $P_{k_{\alpha}}$ in the common eigenbasis of the $P_{k_{\alpha}}$’s. The outcomes are averaged to get an estimate of $\langle\Psi_{V}|P_{k_{\alpha}}|\Psi_{V}\rangle$. The statistical error stemming from the finite number $N_{s}$ of repetitions, or shots, is of the form $\epsilon\approx\sigma(P_{k_{\alpha}})/\sqrt{N_{s}}$ where $\sigma(P_{k_{\alpha}})$ is the standard deviation of individual measurements. Therefore, the run time of computing $E_{V}$, which is proportional to $N_{s}$, scales as $1/\epsilon^{2}$. This scaling is a direct consequence of the fact that energy measurements in VQE are based on a classical averaging of individual measurement outcomes. Even more advanced methods like classical shadows [22] also suffer from this scaling, although they can alleviate the dependence of the run time on the number of terms to be measured.
This unfavorable scaling of VQE leads to quite large run time estimates if good enough accuracies $\epsilon$ are to be reached: in 2015, a study estimated that to evaluate the ground state energy of Fe2Se2 in the minimal STO3-3g orbital basis with chemical precision would require a prohibitive number of $10^{19}$ samples [15]. Another study estimated the runtime to estimate the combustion energy for methane to be of 1.9 days and 71 days for ethanol, finding no practical advantage compared to state-of-the-art classical techniques (see Table II in [14]). In a recent review, combining the state of the art methods in each part of the VQE pipeline and assuming a gate time of $100$ ns, the total runtime for one iteration of VQE on the Cr2 molecule is estimated at 25 days. Taking into account the number of iterations required to reach the minimum in parameter space yields an estimate of 24 years [10].
@@ I.2 The decoherence versus precision criterion
In this section, we disregard statistical uncertainty and barren plateaus, and aim at answering a more pragmatic short-term question: what is the level of decoherence sustainable to attain a targeted precision on the ground state energy? (problem III). Note that we use the terminology ”decoherence” generically here to account for all the imperfections of the quantum hardware including actual decoherence, but also more mundane limitations to the precision such as the effect of crosstalk, $1/f$ noise, imprecise pulse sequence [27].
@@ I.2.1 Derivation of the criterion
We measure the effect of noise on the hardware with the fidelity $F=\langle\Psi_{V}|\rho|\Psi_{V}\rangle$. It expresses how the density matrix $\rho$ of the quantum computer after a noisy execution of the variational circuit $U(\vec{\theta})$ differs from the expected one, $|\Psi_{V}\rangle\langle\Psi_{V}|$. $F<1$ implies that $\rho$ can always be written in the form
@@
with $E_{\rm noise}={\rm Tr}(\rho_{\rm noise}H)$. A large corpus of experiments and theory [28, 29], including the seminal ”quantum supremacy” experiment by Google [30], shows that the fidelity decays exponentially with the total number of applied gates $N_{g}$,
@@
where $\epsilon$ is the average error per gate. This accumulation of errors puts a severe constraint on the number of gates and thereby on the performance of noisy quantum computers [31, 32]. The fact that different errors get combined in a multiplicative way can be put on firm mathematical basis only for certain error models, but is widely expected to hold more generally.
Suppose that one targets a precision $\eta_{\rm chem}$ in the calculation with $\eta_{\rm chem}\ll E_{\mathrm{noise}}-E_{V}$. It follows that $F$ must be very close to unity. Combining the above equations, imposing $\Delta E\leq\eta_{\rm chem}$ leads to $\epsilon<\epsilon_{\rm max}$ with
$\epsilon_{\rm max}=\frac{\eta_{\rm chem}}{(E_{\rm noise}-E_{V})N_{g}}.$ (8)
This is our quantitative criterion for using VQE on a given hardware, molecule and ansatz. Let us emphasize that, in practice, $N_{g}$ depends implicitly on $\eta_{\rm chem}$: good precision requires a complex ansatz, and hence a large number of gates. Note that $\epsilon_{\rm max}$ is the effective error that one obtains after one has done one’s best to decrease the error level, including any quantum error correction scheme and/or error mitigation. Alternatively, one can write the same criterion in terms of the bare error level $\epsilon_{\rm bare}$, i.e. the error level as measured in randomized benchmarking on single qubits and pairs of qubits. In the leading quantum hardware, the average error per gate is dominated by the two-qubit gates for which $\epsilon_{\rm bare}\leq 1\%$. If one uses (possibly costly) error mitigation techniques that reduces the error in post processing by a factor $1/G$, we can write the above estimate in terms of the bare error level by simply replacing,
$1-F\rightarrow G(1-F)$ (9)
which in the limit of small $\epsilon$ ($F$ close to unity) amounts effectively to using an average error per gate given by $\epsilon=\epsilon_{\rm bare}G$. In terms of the bare error, the criterion (8) is therefore slightly less stringent. One arrives at $\epsilon_{\rm bare}<\epsilon_{\rm max,b}$ with
$\epsilon_{\rm max,b}=\frac{\eta_{\rm chem}}{(E_{\rm noise}-E_{V})N_{g}G}.$ (10)
Values as small as G=0.0010.01G=0.001-0.01 have been reported in the literature for relatively small systems, but it is unclear if such large gains can be maintained when the system is scaled up owing to the generically exponential cost of error mitigation (as discussed in more detail in Sec. I.2.6). Since our goal is to find constraints on the hardware and the value of $G$ is strongly hardware dependent, in the following, we focus on the criterion for the effective error level $\epsilon_{\rm max}$.
Figure 1: Schematic of the difference between the hardware spectrum and that of the studied molecule. In terms of the target eigenstates, the VQE ansatz is close to the ground state of the molecule; in terms of the hardware eigenstates, it is made of arbitrary, both low- and high-energy states. On the other hand, the hardware noise can populate arbitrarily high excited states of the studied molecule. For instance, relaxation can populate the hardware ground state, which consists of arbitrarily high energy states in terms of the target eigenstates.
@@ I.2.2 Discussion: scaling of the $(E_{\mathrm{noise}}-E_{V})$ energy scale
We now argue that the energy scale in the denominator, $E_{\mathrm{noise}}-E_{V}$, is generically a large energy scale—of the order of the bare matrix elements of the Hamiltonian, Hartrees or tens of Hartrees—and even generically scales very unfavorably as the square of the number of electrons.
To understand the origin of this behavior, we need to remember that there are two very different Hamiltonians in the problem with two very different energy spectra: the target Hamiltonian $H$ (the one we want to find the ground state of) and the Hamiltonian $H^{\rm hw}$ that describes the quantum hardware (sometimes called the ”resource Hamiltonian”, see e.g., Ref [33]). More generally, the dynamics of the quantum computer hardware is described by a time-dependent Lindbladian, but for the purpose of this discussion, it is sufficient to consider its Hamiltonian. We thus have two sets of eigenstates and eigenenergies,
@@
Our main observation is that in general the target Hamiltonian $H$ is very different from the hardware Hamiltonian $H^{\rm hw}$. Indeed the same hardware Hamiltonian must be used for calculating different molecules so it generically has no correlations with the latter. This is a constraint that a gate-based quantum circuit approach (as opposed to an analog quantum simulation) must obey in exchange for its universality. This situation is different from e.g annealing algorithms where the hardware Hamiltonian is supposed to be slowly deformed into the target Hamiltonian to end up in a low-energy state of the target Hamiltonian. This is also different from analog computing models, which strive to construct hardware Hamiltonians that are as close as possible to the target Hamiltonian, and may thus have similar low-energy subspaces. In other words, gate-based models, to gain universality, lose the physical constraints (for instance, symmetries) that could guarantee that the hardware Hamiltonian indeed shares properties with the target Hamiltonian.
@@
with $c_{i}$ having non zero components in the entire spectrum $E_{i}^{\rm hw}$ (see Fig. 1 for a schematic). Conversely the hardware eigenstates involve very high energies of the molecule (possibly even states with different number of electrons). Why is this observation important? Because the mechanisms that lead to imprecisions and decoherence usually have a simple structure related to the hardware spectrum. This in turn, means that these mechanisms will be structureless with respect to the target spectrum. A simple example is energy relaxation of the hardware—a ubiquitous process. Its fixed point is the thermal density matrix $\rho\propto\sum_{i}e^{-E_{i}/kT}|\Psi^{\rm hw}_{i}\rangle\langle\Psi^{\rm hw}_{i}|$. As another example, consider the admixture of a small amount (1F)(1-F) of, say the ground state, $|\Psi^{\rm hw}_{0}\rangle\langle\Psi^{\rm hw}_{0}|$ in the density matrix: it will lead to a noise energy $E_{\rm noise}=\langle\Psi^{\rm hw}_{0}|H|\Psi^{\rm hw}_{0}\rangle$. Since $|\Psi^{\rm hw}_{0}\rangle$ is composed of a superposition of ”target” states $|\Psi_{i}\rangle$ of arbitrary ”target” energies $E_{i}$, it follows that $E_{\rm noise}$ typically lies in the middle of the spectrum of the target Hamiltonian,
@@
where the magnitude of the constants $a$ and $b$ will depend on the details of the problem and of the noise model. $a$ and $b$ are large energy scales, of the same order as the matrix elements $h_{ij}$ and $v_{ijkl}$ (from Hartrees to tens of Hartrees, depending on the basis set).